OpenStudy (anonymous):

72 is what percent of 480? Write and solve the equation to solve the problem.

OpenStudy (anonymous):

What % of Total Amount is how much The Part. Based off that equation is: % x Total = P Now, fill in the parts: % x 480 = 72 Now for the math portion: p=72/480 p=0.15 0.15 x 100 = 15% Based off that 15% of 480 is 72

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

480 * x = 72 x = 72 / 480 x = 0.15 or 15%

OpenStudy (texaschic101):

percent proportion is : is/of = %/100 72 is x% of 180 72/180 = x/100 cross multiply 180x = (100)(72) 180x = 7200 x = 7200/180 x = 40 check.. 72 = .40(180) 72 = 72 (correct) 40% is your answer

OpenStudy (anonymous):

WAAAAAA???????????????????????????

OpenStudy (texaschic101):

oops...percent formula

OpenStudy (anonymous):

yep its 15% vanessa was right...

OpenStudy (texaschic101):

oops...I wrote it wrong...it is not 180 it is 480 is/of = %/100 72 is x% of 480 72/480 = x/100 cross multiply 480x = 7200 x = 7200/480 x = 15 answer 15%...sorry about the mess up earlier

OpenStudy (anonymous):

I'm confused!

OpenStudy (texaschic101):

you can use the formula : is/of = %/100

OpenStudy (anonymous):

Try typing it rather than numbers. It's easier for me that way.

OpenStudy (texaschic101):

72 is what percent of 480 72 goes in the " in " position what percent means you are looking for the percent, so sub x into the % spot. 480 goes into the " of " position so now we have : 72/480 = x/100 and you just cross multiply and solve for x (the percent)

OpenStudy (texaschic101):

oops...72 goes in the " is " position

OpenStudy (texaschic101):

do you understand or do you need some examples ??

OpenStudy (anonymous):

the x is throwing me off

OpenStudy (texaschic101):

I just substituted a variable into what we were looking for. I substituted x, but you can use any variable.

OpenStudy (anonymous):

I think I need some examples.

undeadknight26 (undeadknight26):

72 divided by 480 gives the percentage boom...

undeadknight26 (undeadknight26):

30/100 is what?

OpenStudy (anonymous):

0.3?

OpenStudy (texaschic101):

15 is 10% of what read it carefully.... is/of = %/100 15 is ...this tells us that 15 goes in the " is " position the 10 goes in the % spot so we are looking for " of " so we sub a variable in the " of " spot giving us : 15/x = 10/100 -- cross multiply 10x = 1500 x = 1500/10 x = 150

undeadknight26 (undeadknight26):

that is the percentage 30 is of 100.

OpenStudy (anonymous):

OOOOOOOOOOOOHHHHHHHHHHHHHHHH!!!!!!!!!!!!!!! Thanks!

OpenStudy (texaschic101):

do you want me to give you a problem to see if you can do it ?

OpenStudy (texaschic101):

what is 10% of 30 formula is : is/of = %/100 can you do this serenity ?

OpenStudy (texaschic101):

we only need 3 number to do this...we have the %, and we have the "of " and we just need to find the " is ".

OpenStudy (anonymous):

I'm good thank you though!

OpenStudy (texaschic101):

ok...I hope you understand it....I hope I was able to help

OpenStudy (anonymous):

You were thanks!

OpenStudy (texaschic101):

good to hear :)